Louise spikes the ball. At the instant the volleyball leaves her hand, its height is 2.6m, and its resultant velocity is 20m/s at an angle of 60 degrees belowhorizontal.

a) How long will it take for the ball to strike the floor if the opposing team does not block it?

b) How far will the ball travel horizontally before it strikes the floor?
Answer #1
a) calculate the speed for the vertical part:
velocity = 20 sin 60 = 17.32 m/s

distance : 2.6 m
velocity : 17.32m/s
gravity: 9.81 m/s^2

2d / t = at + 2(initial velocity)
2(2.6) / t = 9.81t + 2(17.32)
5.2 = 9.81t^2 + 34.64t
solve for t
t = 0.144 or -3.68(not applicable)

Therefore, it will takes 0.144 seconds.

b) find the horizontal speed:
velocity = 20 cos 60 = 10 m/s
distance = 10 * 0.144 = 1.44 m